OpenStudy (anonymous):

log[3]x+log[3]2=2 solve the equation

OpenStudy (anonymous):

log 3(x+2) = 2

OpenStudy (anonymous):

3x+6=2 3x=-4 x=-4/3 ?

OpenStudy (anonymous):

\[\log_3(2x) = 2\]\[3^{2} = 2x\]x = 9/2.

OpenStudy (anonymous):

Ahh, okay. Thanks!

OpenStudy (anonymous):

:^)

OpenStudy (anonymous):

I'm pretty sure that's log 3 (x+2) = 2

OpenStudy (anonymous):

Fortunately not, think about what would happen if logarithms were linear.

OpenStudy (anonymous):

Property of logarithms:\[\log_a(u) + \log_a (v) = \log_a (u v).\]